# Interaction of water surface waves with periodic and quasiperiodic cylinder arrays

### v1

@@ IV.2 Blocking
A numerical summary of these results is compiled in Table 2, where we show the maximum blocking % for any frequency, and the total blocking observed over the full frequency range. Earlier work by Ha et al. using the multiple-scattering method – in which diffracted waves are linked to the incoming wave and represented by Fourier–Bessel expansions to derive the band structure of similar arrays – shows the existence of complete band gaps at for a square array and for an hexagonal array [14]. Our results show blocking dips starting at lower frequencies and bounded by these values at the upper end.
@@
Arrays 4f and 6f, which have quasiperiodicity defined by vacancy defects, as opposed to intrinsic geometry, show effective broadband blocking, duplicating the single broad dip exhibited by 4f and 6f in addition to increased blocking at higher frequencies.
For the quasiperiodic arrays based on or higher rotational symmetries, the blocking curves are characterized by more dips, that individually are narrower than those for the periodic arrays. The behaviour of 4fτ is strikingly similar to 8f, in terms of the locations of the blocking dips. The action of the arrays over a wide angular range indicates that quasiperiodic lattices could contribute to broadband blocking strategies in addition to their rotational isotropy (which we do not investigate here).
@@
IV.3 Resonance
Transmission through the array depends on resonance between some characteristic length in the array and the surface wavelength. Various resonances were observed across the arrays, corresponding with effective transmission. Some examples are shown in Figure 6, in which the intensity maps for the 4f array are reproduced for certain values of . Resonances are indicated, as are datasets far from resonance. These follow a notation in which the numerator is the number of lattice parameters and the denominator is the number of wavelengths. As also shown, when far from resonance, reflection of the waves dominates, forming strong standing wave patterns on the leading side of the arrays and a concomitant shadow on the trailing side.
@@ IV.3 Resonance
The angular frequency is an input variable, and runs from 3.0 – 8.0 rad s-1 in increments of 0.1. The incident wavelength is calculated from . The wavelength inside the array is given by , where is the refractive index calculated from the filling fraction of the array [14]. The actual wavelength is expressed via the resonance ratio, given that the lattice parameter of the array is 1 m.
The difference between actual resonant wavelength and the expected wavelength inside the array is used to calculate the fractional change of wavelength , per wave, that is needed to achieve resonance with the array:
@@
For array 4f, resonance is strongly correlated with transmission, and, to achieve resonance, the wavelength inside the array may be distorted by up to around . Higher order resonances are more distorted in our data, but that may be due to the angular frequency resolution, as the 0.1 rad s-1 steps become more significant at higher angular frequencies. The large degree of distortion achieved to gain resonance is strongly suggestive that the waves will follow such a resonance even if the array undergoes a moderate transformation. We leave this possibility for future work.
@@
The corresponding intensity maps for array 4f, based on a square lattice with elements removed in a Fibonacci sequence, are shown in the lower half of Figure 6 for comparison. The resonances observed in 4f are strongly disrupted in 4f, resulting in reduced transmission. This effect becomes more pronounced with increasing , resulting in the increased blocking of array 4f over the entire frequency range.
V Link to transformation optics
As mentioned in the Introduction, this work has been partly inspired by the rise in transformation-based approaches to wave physics. Generally, this involves transformation (modification of) the constants in a wave propagation equation; permittivity and permeability for optics, Young’s modulus for acoustics, and gravity and depth for water surface waves. Modifying the density of a lattice of elements of given permeability/permittivity [Young’s modulus] causes optical [acoustic] waves to propagate in the direction of the density modifications. The situation for water surface waves is essentially the same. Some wave equations are summarized in Table 3.
Here we have presented the results of linear wave propagation through several array types. We have shown the existence of band gaps for all the arrays, strongly indicating that this is a viable strategy for blocking and/or reflecting wave energy. However, the relevance of our results in the context of transformation optics comes rather from the results shown in Figure 6; specifically, the resonance-led propagation modes. These provide insight into the potential engineering of resonant array waveguides to guide propagation of, rather than block, water surface waves. From simple inspection of Figure 6, it is clear that the waves are in some kind of resonance for a greater range of frequencies than is blocked by band gaps.
optics | acoustics | water surface waves |
VI Conclusions
@@
We have used the open-source Capytaine software to investigate the behaviour of water surface wave array waveguides for a range of array geometries, with the number of array elements ranging from 199-288. Using this approach, we have successfully replicated an earlier result that finds band gaps for some of these array types [14]. Our methodology allows us to directly observe many wave phenomena in real space, for example Bragg diffraction, refraction and resonance. Significantly, our use of simple archetypes of periodic and quasiperiodic lattices demonstrates the potential of this approach to investigate any kind of array waveguide within the linear water wave regime.
@@
Interestingly, transmission through the waveguides is heavily influenced by resonances between the waves and the array geometry. These resonances can be easily disrupted by, e.g., removing array elements, further enhancing the ability of array waveguides to control wave propagation. The resonances suggest that a transformation-optics-inspired approach of modifying the arrays could be harnessed to aid propagation in preferred directions.
Of the arrays tested, the periodic arrays and two quasiperiodic arrays are characterised by a single lattice parameter. The blocking curves of these arrays are characterised by two large blocking dips.
@@
The array with the most effective blocking over the frequency range is hexagonal with quasiperiodically located defects. This array also has the lowest filling fraction of those studied.
@@
Finally, the use of full non-linear potential flow calculations in future applications of this approach would allow investigations of energy transfer between different modes.
VII Acknowledgments
JAS would like to acknowledge useful discussions with P. Moriarty, R. Lifshitz, M. Ancellin and B. A. Patterson. The figures in this manuscript were produced using Inkscape, Plot2 and Veusz. The data were analysed using MATLAB and Fiji.

### v2

@@ 4.1 Blocking
A numerical summary of these results is compiled in Table 2, where we show the maximum blocking % for any frequency, and the total blocking observed over the full frequency range. Earlier work by Ha et al. using the multiple-scattering method – in which diffracted waves are linked to the incoming wave and represented by Fourier–Bessel expansions to derive the band structure of similar arrays – shows the existence of complete Bloch band gaps at for a square array and for a hexagonal array [34]. Our results show almost 100% blocking dips starting at lower frequencies and bounded by these values at the upper end.
@@
In contrast, arrays 4f and 6f, which are identical to 4f and 6f, but with elements removed in a quasiperiodic fashion, have slightly lower blocking in their respective Bloch band gaps. However, they do show improved blocking in regions away from the major blocking dips. This leads to overall better blocking performance over the entire frequency range, as tabulated in Table 2.
Previous analytical approaches to two-dimensional problems have shown that the existence of zero-reflection wave modes is expected if and only if the scattering potential is symmetrical with respect to the wave propagation direction [25, 29]. The behaviour of two-dimensional wave-obstacle systems cannot be meaningfully extrapolated to fully describe three-dimensional interactions between water waves and an array waveguide. However, the lifting of symmetry via the addition of vacancy defects will contribute to the removal of zero-reflection modes, and therefore the increased blocking outside of the band gap. This loss of symmetry also disrupts the primary Bragg resonance, resulting in some transmission of frequencies inside the band gap.
A hexagonal lattice is denser than a square lattice of the same lattice parameter. Therefore, of the periodic arrays, 6f has the highest filling fraction, which seems a reasonable explanation of the slightly better blocking performance of the hexagonal arrays. When this is combined with symmetry removal via the addition of vacancy defects, the combination of low filling fraction and optimal blocking performance is achieved.
For the quasiperiodic arrays based on , or on higher rotational symmetries, the blocking curves are characterized by more dips, that individually are narrower than those for the periodic arrays. The behaviour of 4fτ is strikingly similar to 8f, in terms of the locations of the blocking dips. The action of the arrays over a wide frequency range indicates that quasiperiodic lattices could contribute to broadband blocking strategies in addition to their rotational isotropy (which we do not investigate here). The quasiperiodic arrays cannot by definition be perfectly periodic, or symmetrical, in the propagation direction, which limits the possibilities for both zero-reflection modes and a Bloch band gap.
@@
4.2 Bragg diffraction
In x-ray and electron crystallography, a beam of x-rays or electrons is made incident on a crystal. The scatterers (atoms) in a crystal can be grouped into families of parallel planes of atoms, where one is differentiated from another via its Miller indices. The Miller indices are the numbers of unit cells in each direction to define a family of planes via its normal. For example, (001) in a cubic crystal involves moving 0 unit cells in , 0 unit cells in and 1 unit cell in ; this defines the normal to, and thus refers to, the -planes [(010) refers to the -planes, (100) refers to the -planes]. Each unique family of planes has a unique set of Miller indices, and produces a family of harmonic peaks in the diffraction pattern. The angle at which a peak is located is given by the Bragg law , where is the separation between planes.
In a two-dimensional array of scatterers, the planes are now lines of scatterers, with particular separations. In general, all of the arrays produce strong diffraction when the Bragg condition is satisfied by one or more separations in the array. The diffracted beams are directional and diffuse, which is consistent with the small number of scatterers in the arrays[23].
Here, we use the Bragg law to extract the apparent interlinear distance of the arrays when the wavefront is normal to the array. For infinite periodic arrays is a simple function of , the lattice parameter. For finite periodic arrays, it is a function of with an error due to the finite radius of the scatterers.
For quasiperiodic arrays, there is no single lattice parameter, though examination of the arrays reveals that their geometries depend on arrangements of two characteristic separations. For example, in the Penrose lattice 10f, these separations are 1 m, the rhombus edge length, and m, the width of narrow rhombi. Diffraction peaks then arise from every possible combination of these distances, with relative intensities given by the relative density of occurrences of each combination. For more discussion on this topic, the reader is referred to e.g. [61, 51].
All the arrays have been simulated over the range . The procedure we have used in Figure 5 is:
1.
Identify the -fold symmetry of the array
2.
Select the value of that produces diffraction at the expected angle (i.e., through the side of the -gon adjacent to the side impinged upon by incoming waves)
3.
Calculate from known and using the Bragg law
Close to the Bragg angles for a given lattice, constructive interference may occur for a range of frequencies, with the diffracted beam sweeping a concomitant range of angles. Therefore, it is appropriate to use the Bragg law to extract only when a lattice has one or more easily identified directions of symmetry. Figure 5 shows this analysis for the 4f, 6f, 8f, and 10f arrays, where a black arrow indicates the incoming and diffracted wave direction and angle associated with the rotational symmetry of the array (). For the simple square array, the 2D Miller indices of the line family are (11), and is found to be close to the expected value of . For the simple hexagonal array, the lines are in the 2d hcp (100) family. Adjacent lines in this family have ; we observe close to twice this value, as this is the Bragg reflection available in our frequency range. The deviation from the expected values is consistent with the finite size of the array and the finite radius of the scatterers. For example, for array 4f, scattering is evident from a continuous range of lattices between those defined by the innermost points and outermost points of the cylinders in the array. In any finite periodic array, the number of elements and the dimensions of the array lead to a minimum total number of lines () belonging to any set of Miller indices; in array 4f, at an angle of rad, (Miller indices (11)), is 11. This gives an approximate % error from the limits of , given by (, where is the cylinder radius.
The 4f, 6f, and 6f arrays show the same broad behaviour as the simple structures with this analysis. The 8f and 10f arrays have well-defined scattering angles, which permits the extraction of an apparent ‘interlinear’ distance, shown in the figure as .
The -scaled 4–fold array 4fτ, not shown here, has ambiguous behaviour: although it has 4–fold symmetry, it never strongly diffracts from a set of lines at an angle of rad as expected. However, weak diffraction at multiple wavelengths yields values for of 1.00 m and 1.39 m.
4.3 Lattice coherences
In our data, transmission of surface wave energy through an array (visualised as intensity downstream of the array) is usually observed simultaneously with a periodic variation in intensity inside the array. The period is expressed via a simple ratio between the array periodicity and the incident surface wavelength. We call this periodic variation in intensity a lattice coherence (LC).
Various LCs were observed across the arrays, corresponding with effective transmission. In each case, the wavelength is distorted from the incident wavelength to fit the LC. The pattern of intensity within the array is examined and the period and hence unit cell determined via feature correlation. Following this, the number of wave crests in the unit cell of the LC is counted. The number of lattice periods is readily countable by superimposing the array on the data.
In Figure 6, we explore the intensity maps for the 4f array, the simplest case, and the structurally similar 4f array, for certain values of . The angular frequency is an input variable, and runs from 3.0 – 8.0 rad s-1 in increments of 0.1. In each panel for the 4f array, we list certain values:
: the angular frequency of the incident surface wave (rad s-1)
: a wave-lattice structure interaction parameter, provided for comparison to the work of Ha et al.[34]
: the incident surface wavelength, calculated from
@@ 4.3 Lattice coherences
: the wavelength of the wave inside the array adjusted for the lattice index of refraction, given by , where is the refractive index calculated from the filling fraction of the array [34]
: the fractional difference between and the LC ratio, or wavelength inside the array, ; this is a measure of the distortion the waves undergo to reach LC
LC ratio (at the bottom of each relevant panel): the ratio of number of lattice periods () to number of wavelengths (); this is numerically equal to the wavelength of the waves inside the array, given that the lattice period is 1 m.
Identifiable LCs are indicated; these follow a notation in which the numerator is the number of lattice parameters and the denominator is the number of wavelengths. A graphical representation of the LC is also shown, showing the relationship between the surface waves (white) and the array periodicity (red).
We borrow a phrase from condensed matter physics: higher order commensurate (HOC), to describe the situation where a LC is visible via a periodic pattern of intensity, but where a complete cycle of the LC does not fit inside the array.
For array 4f, LCs are strongly correlated with transmission, and, to achieve LC, the wavelength inside the array may be distorted by up to in our data. A LC results in a rational relationship between wave and array: waves at the same position in each unit cell of a LC encounter a scatterer at the same angle of phase. Each unit cell is therefore symmetrical; the transmission function for each unit cell therefore has zero-reflection modes, and the stacking of unit cells across the array extends the zero-reflection modes across the array.
@@
The large degree of distortion achieved to gain LC is strongly suggestive that the waves will follow such a LC even if the array undergoes a moderate transformation. We leave this possibility for future work.
@@
The corresponding intensity maps for array 4f, based on a square lattice with elements removed in a Fibonacci sequence, are shown in the lower half of Figure 6 for comparison. In analogy to the scattering of electron waves by defects in periodic crystals [30], the LCs observed in 4f are strongly disrupted by the defects in 4f, resulting in reduced transmission. The mechanism for this disruption is the removal of symmetry and periodicity, and therefore zero-reflection modes.
5 Conclusions
@@
We have used the open-source Capytaine software to investigate the behaviour of water surface wave array waveguides for a range of array geometries, with the number of array elements ranging from 199-288. Using this approach, we have successfully replicated an earlier result that finds band gaps for some of these array types [34], and indeed have shown the existence of band gaps for all the arrays, strongly indicating that this is a viable strategy for blocking and/or reflecting wave energy. Our methodology allows us to directly observe many wave phenomena in real space, for example Bragg diffraction, refraction and resonance. Significantly, our use of simple archetypes of periodic and quasiperiodic lattices demonstrates the potential of this approach to investigate any kind of array waveguide within the linear water wave regime.
@@
Of the arrays tested, the periodic arrays and two quasiperiodic arrays are characterised by a single lattice parameter. The blocking curves of these arrays are characterised by two large blocking dips caused by the primary Bragg resonance and formation of a Bloch band gap.
@@
The array with the most effective blocking over the frequency range is hexagonal with quasiperiodically located vacancy defects. This array also has the lowest filling fraction of those studied.
Transmission through the waveguides is heavily influenced by relationships between the wavelength and the array geometry, here called lattice coherences (LC). These LCs can be easily disrupted by, e.g., removing array elements, further enhancing the ability of array waveguides to block wave propagation. Here, we have placed vacancy defects in a systematic quasiperiodic fashion. Future work could explore different geometries of vacancies, for example, periodic or random.
Beyond this, the observation, from simple inspection of Figure 6, that the waves are in some kind of LC for a greater range of frequencies than is blocked by band gaps, strongly suggests that it may be more effective to control waves by using the transformation-optics-inspired approach of modifying the arrays to aid propagation in preferred directions.
@@
The use of full non-linear potential flow calculations in future applications of this approach would allow investigations of energy transfer between different modes.
Finally, the results highlight the need for experimental investigations of array waveguides. Here, we probe only the linear regime, and find significant opportunities. When the full range of water surface wave phenomena are allowed to interact with array waveguides, additional avenues for research may become apparent.
6 Acknowledgments
JAS would like to acknowledge useful discussions with P. Moriarty, R. Lifshitz, M. Ancellin and B. A. Patterson. The figures in this manuscript were produced using Inkscape, Plot2 and Veusz. The data were analysed using MATLAB and Fiji (ImageJ).